OpenStudy (anonymous):

simplify the square root of 20y^8

OpenStudy (anonymous):

can u draw it plz

iYuko (iyuko):



Nnesha (nnesha):

\[\huge\rm \sqrt{20y^8}\] factro 20 and change root to exponent exponent rule!\[ \huge\rm \sqrt[n]{x^m}= x^\frac{ m}{ n}\]

Nnesha (nnesha):

factor*